# Balancing Embedding Spectrum for Recommendation

### v1

@@ III-B Incomplete Collapse
where $\mathbf{\bar{A}}_{uj}=\sigma(\hat{r}_{uj})$ for $(u,j)\in\mathbf{R}^{-}$. As the direction of the gradient is opposite to that of observed interactions, it seems that the disconnected nodes (unobserved interactions) are pushed away from the target users/items, preventing the representations from collapsing.
However, we now show that negative sampling alone is actually not able to avoid the complete collapse. Ideally, with the magnitude of gradient fixed, Equation (9) can be rewritten to the following Equation similar to Equation (5):
$\mathbf{H}^{(l)}=\left(\mathbf{I}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)\right)^{l}\mathbf{H}^{(0)}.$ (10)
@@
Here, $\mathbf{I}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)$ can be disentangled to a low pass $\mathbf{I}+\alpha\mathbf{A}$ and a high pass filter $\mathbf{I}-\alpha\mathbf{\bar{A}}$, where a high pass filter can balance the embeddings spectrum by reducing the weights on low frequency and raising the importance of high frequency components, leading to a more uniform distribution. We visualize the spectrum of $\mathbf{A}-\mathbf{\bar{A}}$ in Figure 3 (a), where we fix the elements of $\mathbf{A}$ and $\mathbf{\bar{A}}$ to 1 and normalize them separately. $\mathbf{\bar{A}}=0$ (zero matrix) for ‘Only positive’, and $\mathbf{\bar{A}}$ contains all unobserved interactions for ‘With negative’. It is obvious that $\mathbf{A}-\mathbf{\bar{A}}$ has a more balanced spectrum than $\mathbf{A}$, leading to a higher erank. However, according to our previous analysis in Section III-A, the largest eigenvalue still would dominate while other smaller eigenvalues are filtered out as updating the algorithm enough times (i.e., $\frac{(1+\alpha\lambda_{k})^{l}}{(1+\alpha\lambda_{1})^{l}}\rightarrow 0$ ($k\neq 1$) ) as long as the spectrum of $\mathbf{I}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)$ is not completely uniform, resulting in a complete collapse. To verify our analysis, we adopt Euclidean distance loss whose magnitude of gradients does not change with training, and show how the erank changes on the model with only positive and both positive and negative samples in Figure 3 (b). We can observe that the erank of the model exploiting negative samples still monotonically declines, regardless of a slower speed than the model with only positive samples. Then, another question rises: What enables BCE and BPR loss to avoid the complete collapse?
As shown in Figure 2 (c), the term controlling the magnitude of the gradient over the positive sample $\mathbf{A}_{ui}$: $1-\sigma(\hat{r}_{ui})$ is way larger than 0 at the beginning of training, causing the embeddings to collapse. Since positive pairs are more likely to be sampled due to the data sparsity, the gradients vanish faster as they are pulled close: $\mathbf{A}\rightarrow 0$, allowing the high pass filter to balance the spectrum. Then, negative samples being pushed away also leads to $\mathbf{\bar{A}}\rightarrow 0$, making the representations converge: $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}$ and leading to an incomplete collapse. Furthermore, since it has been shown that raising the negative samples leads to a desirable representation and improvement [5, 13], we evaluate BCE with different negative sampling ratios and report the results in Figure 3 (c). We can see that the accuracy increases by raising the sampling ratio, while the erank remains unchanged, showing that introducing more negative samples fails to further alleviate or prevent the embedding collapse. To conclude, gradient vanishing on log loss (including BCE and BPR loss) is the ‘X factor’ to prevent the complete collapse, and is also the reason causing the incomplete collapse as well. Therefore, a more effective design proved to completely solve the collapse issue is required.
TABLE I: A toy example demonstrating the effectiveness of Algorithm 1. The matrix is randomly generated with a size 10. Without specification, $T=50$, $\alpha=1e-3$, and $L=1$.
$T$ 0 1 10 100 500 1000
erank 7.64 7.72 8.11 8.59 9.57 9.75
$\alpha$ 0 1e-3 2e-3 5e-3 1e-2 2e-2
erank 7.64 8.43 8.71 9.09 9.35 9.59
$L$ 0 1 2 3 - -
erank 7.64 7.80 8.45 8.83 - -
IV Methodology
IV-A Balancing Embedding Spectrum
The spectrum distribution can directly reflect the extent of embedding collapse. Intuitively, a uniform distribution indicates that different dimensions equally contribute to the embeddings and leads to a high effective rank. Therefore, a reliable solution is an all pass filter $\mathcal{F}^{A}$ such that $\mathcal{F}^{A}(\sigma_{i})$=$\mathcal{F}^{A}(\sigma_{j})$ for $\sigma_{i}\neq\sigma_{j}$. By parameterizing $\mathcal{F}^{A}$, it is straightforward to directly balance the spectrum by optimizing the following loss:
${\rm min}\sum_{k\neq z}\left\|\sigma^{\prime}_{k}-\sigma^{\prime}_{z}\right\|^{2}_{2},$ (11)
where $\sigma^{\prime}_{k}$ is the singular value of $\mathcal{F}^{A}\left(\mathbf{H}\right)$. However, such a design requires the computation of singular values during each training and introduces more model parameters that might hinder the model convergence, thus we stick to a simple yet effective design in this work. Consider an affinity graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathbf{\hat{A}})$, where the node features are represented by the embedding matrix $\mathbf{H}$ and the adjacency relation is measured by the similarity score between nodes: $\mathbf{\hat{A}}_{ui}=\mathbf{h}_{u}^{T}\mathbf{h}_{i}$, we define the message passing on $\mathcal{G}^{\prime}$ as follows:
$\mathcal{F}^{A}\left(\mathbf{H}\right)=\mathbf{H}-\alpha\mathbf{\hat{A}}^{L}\mathbf{H},$ (12)
where $\alpha\in\mathbb{R}^{+}$, $L\in\mathbb{N}$ is the matrix power.
@@ IV-A Balancing Embedding Spectrum
$\overbrace{\mathcal{F}^{A}\cdots\mathcal{F}^{A}}^{T}(\mathbf{H})$ (Repeating $\mathcal{F}^{A}(\mathbf{H})$ $T$ times ) is equivalent to an all pass filter where ${\rm erank}(\overbrace{\mathcal{F}^{A}\cdots\mathcal{F}^{A}}^{T}(\mathbf{H}))=d$
Input : Embedding matrix $\mathbf{H}$; matrix power $L$; the number of iteration $T$; scaling hyperparameter $\alpha\in\mathbb{R}^{+}$
for $t=1$ to $T$ do
      $\mathbf{H}\leftarrow\mathbf{H}-\alpha{\mathbf{\hat{A}}}^{L}\mathbf{H}$;
      end for
Return $\mathbf{H}$
Algorithm 1 Balancing Embedding Spectrum via Message Passing on $\mathcal{G}^{\prime}$
Input : Embedding matrix $\mathbf{H}$; batch size $B$; scaling hyperparameter $\alpha\in\mathbb{R}^{+}$
sample the users and items: $\mathcal{U}_{B}$, $\mathcal{I}_{B}$ ;
@@
Return $\mathbf{H}_{B}^{U}$, $\mathbf{H}_{B}^{I}$
@@
We can rewrite Equation (12) according to singular value decomposition (SVD) as follows:
@@
$\mathcal{F}^{A}\left(\mathbf{H}\right)$ $=\mathbf{H}-\alpha\left(\mathbf{H}\mathbf{H}^{T}\right)^{L}\mathbf{H}$ (13)
$=\mathbf{P}diag\left(\sigma_{k}\right)\mathbf{Q}^{T}-\alpha\mathbf{P}diag\left(\sigma_{k}^{2L+1}\right)\mathbf{Q}^{T}$
$=\mathbf{P}diag\left(\sigma_{k}\left(1-\alpha\sigma^{2L}_{k}\right)\right)\mathbf{Q}^{T},$
where $\mathbf{P}$ and $\mathbf{Q}$ are stacked singular vectors, $diag(\cdot)$ is the diagonalization operation. It is obvious that $1-\alpha\sigma^{2L}_{i}>1-\alpha\sigma^{2L}_{j}$ for $\sigma_{i}<\sigma_{j}$. Thus $1-\alpha\sigma_{k}^{2L}$ can be considered as rescaled factors: the larger singular values corresponding to the lower frequency components are multiplied by smaller weights, leading to a more uniform distribution when $L>0$. $L$ and $\alpha$ control the weight multiplied on the singular value of $\mathbf{H}$, thus directly affect the efficiency of spectrum balancing. A larger (smaller) $L$ makes the rescaled factors smaller (larger) on the large singular values, and $\alpha$ controls the norms of the rescaled factors that has a similar effect to $L$. Since $\sigma_{k}^{L}$ increases fast, a large $L$ could reduce the algorithm’s efficiency where the large singular values in the original distribution are multiplied by too small weights, causing a sharp singular value distribution similar to the original one. In addition, too large $L$ or $\alpha$ could also break the non-negativity of the singular value, thus we need to assure that $01-\alpha\sigma_{1}^{2L}>$ 0. According to the definition of Entropy and erank, a more uniform distribution results in a larger erank, indicating that $erank(\mathcal{F}^{A}(\mathbf{H}))\geq erank(\mathbf{H})$ where $=$ holds when $erank(\mathbf{H})=d$. By sufficiently repeating Equation (13), eventually $erank(\mathbf{H})=d$. We summarize the above analysis and propose Algorithm 1 to tackle the embedding collapse. We run Algorithm 1 on a randomly generated matrix with the code ${\rm torch.randn(10,10)}$ [^1] https://pytorch.org/, and report the results with different settings of $\alpha$, $L$, and $T$ in Table LABEL:algorithm_1_effective, where $01-\alpha\sigma_{1}^{2L}<$ 0 when $L>3$. We can see that the erank tends to converge to $d=10$ as we increase the iteration $T$, scaling hyperparameter $\alpha$, and the order $L$, demonstrating the effectiveness of Algorithm 1 balancing the embedding spectrum.
Algorithm 1 is impractical since the complexity is $\mathcal{O}(TLD^{2}d)$ where $D=|\mathcal{U}|+|\mathcal{I}|$. We propose a simplified implementation of DirectSpec formulated in Algorithm 2. Since the singular value increases fast with $L$ which could break the non-negativity of the singular value, and in practice we do not need the erank to strictly be $d$ (we will explain in Section IV-C), we can set $T=1$ and $L=1$. In addition, instead of updating the whole embedding matrix, we iteratively rebalance the spectrum of a smaller matrix with a size in accordance with the batch size.
Figure 4: An illustration of the proposed DirectSpec.
IV-B A Decorrelation Perspective
In this subsection, we show how users and items are represented in the embedding space through our proposed DirectSpec from a spatial perspective.
@@ IV-B A Decorrelation Perspective
Algorithm 2 is equivalent to optimizing the following loss:
$L_{ds}=\sum_{u\in\mathcal{U}}\sum_{v\in\mathcal{U}}\left\|\mathbf{h}_{u}^{T}\mathbf{h}_{v}\right\|^{2}+\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}\left\|\mathbf{h}_{i}^{T}\mathbf{h}_{j}\right\|^{2}.$ (14)
By calculating the gradient over user and item embeddings:
$\frac{\partial L_{ds}}{\mathbf{h}_{u}}=4\sum_{v}\left(\mathbf{h}_{u}^{T}\mathbf{h}_{v}\right)\mathbf{h}_{v},\,\frac{\partial L_{ds}}{\mathbf{h}_{i}}=4\sum_{j}\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}\right)\mathbf{h}_{j},$ (15)
the parameter updating through SGD can be formulated as follows:

### v2

@@ 3.2. Incomplete Collapse
where $\mathbf{\bar{A}}_{uj}=\frac{\partial\mathcal{L}}{\partial\hat{r}_{uj}}$ for $(u,j)\in\mathbf{R}^{-}$. Note that the gradients for positive and negative samples always have the opposite direction. Here, $\mathbf{I}+\alpha\left(\mathbf{A}-\mathbf{\bar{A}}\right)$ can be disentangled to a low pass $\mathbf{I}+\alpha\mathbf{A}$ and a high pass filter $\mathbf{I}-\alpha\mathbf{\bar{A}}$, where a high pass filter can balance the embeddings spectrum by reducing the weights on low frequency and raising the importance of high frequency components, leading to a more uniform distribution. We visualize the spectrum of $\mathbf{A}-\mathbf{\bar{A}}$ in Figure 2 (c), where we fix the elements of $\mathbf{A}$ and $\mathbf{\bar{A}}$ to 1 and normalize them separately. $\mathbf{\bar{A}}=0$ (zero matrix) for ‘Only positive’, and $\mathbf{\bar{A}}$ contains all unobserved interactions for ‘With negative’. It is obvious that $\mathbf{A}-\mathbf{\bar{A}}$ has a more balanced spectrum than $\mathbf{A}$, leading to a higher erank.
@@
Negative sampling alone leads to collapse.
Suppose we only focus on the negative sampling with the updating rule as: $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}-\alpha\mathbf{\bar{A}}\mathbf{H}^{(l)}$. Ideally (i.e., with the gradients remaining unchanged), the updating rule can be rewritten as follows:
(10) $\mathbf{H}^{(l)}=\left(\mathbf{I}-\alpha\mathbf{\bar{A}}\right)^{l}\mathbf{H}^{(0)}.$
Similar to the analysis in Section 3.1, the high frequency components gradually dominate ($\frac{(1-\alpha\lambda_{k})^{l}}{(1-\alpha\lambda_{n})^{l}}\rightarrow 0$) as $l$ becomes larger, and eventually ${\rm rank}\left(\mathbf{H}^{(l)}\right)\rightarrow 1$. Then, another question arises: If both positive and negative sampling lead to collapse, how do existing methods avoid the collapse?
We further zoom in on the gradient over BCE loss. The gradient over the positive sample $\mathbf{A}_{ui}$: $1-\sigma(\hat{r}_{ui})$ is close to 1 at the beginning of training, causing the embeddings to collapse. Since positive pairs are more likely to be sampled due to the data sparsity, the gradients vanish faster as they are pulled close: $\mathbf{A}\rightarrow 0$, allowing the high pass filter to balance the spectrum. As the negative samples are sampled and pushed away, we have: $\mathbf{\bar{A}}_{uj}=\sigma(\hat{r}_{uj})\rightarrow 0$, making the representations converge: $\mathbf{H}^{(l+1)}=\mathbf{H}^{(l)}$ and leading to an incomplete collapse. For BPR loss, since the magnitude of gradients is $1-\sigma(\hat{r}_{ui}-\hat{r}_{uj})$ for both positive and negative pairs, positive pairs being optimized causes the gradients for negative pairs to vanish as well, making the high pass filter have a weaker effect balancing the spectrum and also explaining why BPR(MF) converges more slowly than BCE(MF). To conclude, negative sampling is not designed to prevent the collapse issue, gradient vanishing on log loss is the ‘X factor’ to alleviate the complete collapse, and is also the reason causing the incomplete collapse as well. Therefore, a more effective design proved to completely solve the collapse issue is required.
4. Methodology
4.1. Directly Balancing Embedding Spectrum (DirectSpec)
The spectrum distribution can directly reflect the extent of embedding collapse. Intuitively, a uniform distribution indicates that different dimensions equally contribute to the embeddings and leads to a high effective rank. To this end, we consider the following Lagrangian:
(11) $\mathcal{L}=f\left(\sigma_{1},\cdots,\sigma_{d}\right)+\beta\left(\sqrt{\sum_{i}\sigma_{i}^{2}}-c\right),$
with the minimum reached at $\sigma_{1}=\cdots=\sigma_{d}=\frac{c}{\sqrt{d}}$, where the singular value is bounded by the $L_{2}$ ball with radius $c$.
Proposition 1.
Let $\frac{\partial f}{\partial\sigma_{i}}=g_{i}(\cdot)$. The minimum of Equation (11) is achieved at $\sigma_{1}=\cdots=\sigma_{d}=\frac{c}{\sqrt{d}}$ as long as $g_{i}(\sigma_{j})=g_{j}(\sigma_{j})$ for $i,j\in\{1,\cdots,d\}$.
By calculating the derivative over $\sigma_{i}$, we get:
(12) $\left\{\begin{array}[]{lr}g_{1}(\sigma_{1})+2\beta\sigma_{1}=0\\ \\ \qquad\;\;\cdots\\ \\ g_{d}(\sigma_{d})+2\beta\sigma_{d}=0\\ \\ s.t.\sum_{i}\sigma_{i}^{2}=c^{2}.\\ \end{array}\right.$
Obviously, $\sigma_{1}=\cdots=\sigma_{d}$ as long as $g_{i}(\cdot)$ has the same form for all singular values. For instance, a simple design is $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}$ where $\beta=-\frac{\sqrt{d}}{2c}$.
While the algorithm above has a provable guarantee to prevent the collapse issue, the efficiency is limited as it requires computing the singular values with the complexity as: $\mathcal{O}(\left|\mathcal{U}\right|d^{2}+\left|\mathcal{I}\right|d^{2})$.
Proposition 2.
When $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}^{2K}$, $K\in\mathbb{N}_{+}$ and $K\geq 2$, we can balance the spectrum with the following Equation:
(13) $\mathbf{H}\leftarrow\mathbf{H}-\alpha\left(\mathbf{H}\mathbf{H}^{T}\right)^{K-1}\mathbf{H}.$
Input : Embedding matrix $\mathbf{H}$; batch size $B$; hyperparameter $\alpha\in\mathbb{R}^{+}$
sample the users and items from $\mathbf{R}^{+}$: $\mathcal{U}_{B}$, $\mathcal{I}_{B}$ ;
@@ 4.1. Directly Balancing Embedding Spectrum (DirectSpec)
update parameters via $\mathcal{L}(\hat{r}_{ui},r_{ui}),(u,i)\in\mathbf{R}^{+}$
@@
According to SVD: $\mathbf{H}=\mathbf{P}diag(\sigma_{i})\mathbf{Q}^{T}$ with $\mathbf{P}$, $\mathbf{Q}$, and $diag(\cdot)$ as the left, right stacked singular vectors, and the diagonalization operator, respectively. It is obvious that:
(14) $\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}=\left(\mathbf{P}diag\left(\sigma_{i}^{2}\right)\mathbf{P}^{T}\right)^{K}=\mathbf{P}diag\left(\sigma_{i}^{2K}\right)\mathbf{P}^{T},$
due to the orthogonality of singular vectors: $\mathbf{Q}^{T}\mathbf{Q}=\mathbf{I}$, $\mathbf{P}^{T}\mathbf{P}=\mathbf{I}$. Then, we can make the following observation:
(15) $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}^{2K}=\operatorname{Tr}\left(\mathbf{P}diag\left(\sigma_{i}^{2K}\right)\mathbf{P}^{T}\right)=\operatorname{Tr}\left(\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}\right),$
where the trace $\operatorname{Tr}(\cdot)$ is equal to the sum of eigenvalues. Then, the embedding can be updated via SGD as follows:
@@
(16) $\mathbf{H}\leftarrow\mathbf{H}-\alpha\frac{\partial\operatorname{Tr}\left(\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}\right)}{\partial\mathbf{H}}=\mathbf{H}-\alpha K\left(\mathbf{H}\mathbf{H}^{T}\right)^{K-1}\mathbf{H}.$
By simply setting $\alpha\leftarrow\alpha K$, we obtain Equation (13). Note that we can ignore the constraint term (i.e., $\sum_{i}\sigma_{i}^{2}=c^{2}$) as it can be replaced by the normalization: $\mathbf{H}_{i}\leftarrow\frac{\mathbf{H}_{i}}{\left|\mathbf{H}_{i}\right|}$ where $\mathbf{H}_{i}$ represents an arbitrary row of $\mathbf{H}$, simply due to the following relation:
(17) $\sum_{i}\sigma_{i}^{2}=\operatorname{Tr}\left(\mathbf{P}diag\left(\sigma_{i}^{2}\right)\mathbf{P}^{T}\right)=\operatorname{Tr}(\mathbf{H}\mathbf{H}^{T})=\sum_{i}\mathbf{H}_{i}\mathbf{H}_{i}^{T}=d,$
where the diagonal element $\mathbf{H}_{i}\mathbf{H}_{i}^{T}=1$. To further shed light on how Equation (13) balances the spectrum, we can rewrite it as follows:
(18) $\mathcal{F}\left(\mathbf{H}\right)=\mathbf{H}-\alpha\left(\mathbf{H}\mathbf{H}^{T}\right)^{K}\mathbf{H}$ $=\mathbf{P}diag\left(\sigma_{i}\right)\mathbf{Q}^{T}-\alpha\mathbf{P}diag\left(\sigma_{i}^{2K+1}\right)\mathbf{Q}^{T},$
$=\mathbf{P}diag\left(\sigma_{i}\left(1-\alpha\sigma^{2K}_{i}\right)\right)\mathbf{Q}^{T},$
where we redefine $K\geq 1$ for simplicity. As the function $\sigma_{i}^{2K}$ increases faster than $\sigma_{i}$, $1-\alpha\sigma^{2K}_{i}>1-\alpha\sigma^{2K}_{j}$ for $\sigma_{i}<\sigma_{j}$ as long as $\alpha$ is properly set (i.e., $01-\alpha\sigma_{1}^{2K}>$ 0). Thus, $1-\alpha\sigma_{i}^{2K}$ can be considered as rescaled factors: the larger singular values are multiplied by smaller weights, leading to a balanced distribution. Note that each update would not shrink the singular values as the embedding normalization keeps their $L_{2}$ norm unchanged as shown in Equation (17). According to the definition of entropy and erank, a more uniform distribution results in a higher erank, indicating that $erank(\mathcal{F}\left(\mathbf{H}\right))\geq erank(\mathbf{H})$ where $=$ holds when $erank(\mathbf{H})=d$. To further enhance the efficiency of DirectSpec, we set $K=1$ and adopt a batch training strategy instead of updating the whole $\mathbf{H}$. We can integrate DirectSpec as a spectrum-normalization layer in the recommendation algorithm as formulated in Algorithm 1 and as illustrated in Figure 3 (a), where we do not need to explicitly sample the negative pairs, saving the training cost. To demonstrate and validate the effectiveness of DirectSpec, we run a toy example as shown in Figure 3 (b).
Figure 3. (a): An illustration of the proposed DirectSpec, (b): A toy example running DirectSpec on a randomly generated matrix with torch.randn(10,10) $\{5,10\}$ times.
4.2. Spectrum Balancing Could be Poisonous
@@
The goal of spectrum balancing and alignment of positive pairs is contradictory.
Suppose the user $u$ and $v$ both interacted with item $i$. Let $\mathbf{H}_{U}$ denote the embedding matrix for users, then $\mathbf{H}_{U}\mathbf{H}_{U}^{T}=\mathbf{I}$ when the perfectly balanced spectrum is achieved: $\sigma_{1}=\cdots=\sigma_{d}$, indicating that $u$ and $v$ are orthogonal in the embedding space: $\mathbf{h}_{u}^{T}\mathbf{h}_{v}=0$. On the other hand, $i$ is optimized to be parallel to both $u$ and $v$: $\mathbf{h}_{u}^{T}\mathbf{h}_{i}=1$ and $\mathbf{h}_{v}^{T}\mathbf{h}_{i}=1$ when the perfect alignment is reached, while such a relation is impossible in Euclidean space: a vector is both parallel to two vectors orthogonal to each other. Therefore, balancing spectrum could be poisonous to alignment of positive pairs and vice versa. In figure 4, we investigate how erank, training loss, and accuracy (nDCG@10) change with the intensity of spectrum balancing by adjusting $\alpha$ of DirectSpec. (i.e., larger (smaller) $\alpha$ indicates intenser (weaker) spectrum balancing). We can see that erank rises as increasing $\alpha$, reaching almost $d$ at $\alpha=0.1$. In the meanwhile, the training loss representing the distance among positive pairs (i.e., lower (higher) loss indicates a better (worse) alignment of positive pairs) accordingly rises with erank in Figure 4 (a), demonstrating the contradiction between spectrum balancing and alignment. The accuracy increases and then drops with $\alpha$, reaching the maximum at $\alpha=0.02$ and erank$=62.9$ in Figure 4 (b), showing that the perfect balanced spectrum does not necessarily lead to better performance. In conclusion, a good representation lies in the trade-off between alignment and spectrum balancing, rather than focusing solely on either of them.
Figure 4. How (a): the effective rank and training loss (b): the effective rank and accuracy change with $\alpha$.
4.3. Individualized Spectrum Balancing (DirectSpec+)
As pointed out in Section 4.2 that balancing the embedding spectrum via DirectSpec is equivalent to the orthogonalization of unobserved pairs with each user/item being optimized as:
(19) $\mathbf{h}_{i}\leftarrow\mathbf{h}_{i}-\sum_{j}\alpha\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}\right)\mathbf{h}_{j}.$
Compared to $\alpha$, $\mathbf{h}_{i}^{T}\mathbf{h}_{j}$ can be considered as an individualized coefficient controlling the intensity for balancing the spectrum between $i$ and $j$. Given that the goal is not solely to generate a perfectly spectrum but rather to balance the trade-off between alignment and spectrum optimization, we argue that DirectSpec does not fully consider the following two factors:
  • The intensity should be adaptive to the extent of collapse (i.e., the pairs suffering more (less) from the collapse issue should be balanced with a higher (lower) intensity ).
  • Unobserved pairs may also show relations with varying degrees (i.e., the pairs showing closer relations should be balanced with a lower intensity).
We propose an enhanced individualized spectrum balancing design DirectSpec+ as follows: